OpenStudy (error1603):

Determine f'(x) for f(x) = x22 - xy + sin x cos x - ex

sam (.sam.):

\[f(x) = x^{22} - xy + \sin x \cos x - e^x\] You're asked to find \(f'(x)\)?

OpenStudy (error1603):

Differentiate f(x) with respect to x?

OpenStudy (mrnood):

@Error1603 how does this relate to your LAST question about slices of cake - they are in a different league in terms of level of maths....

OpenStudy (error1603):

So, f'(x) = 2x - y + (sin x(-sin x) + cos x ( cos x) ) - ex

OpenStudy (error1603):

Im stuck now

OpenStudy (error1603):

Welp I guess I get no help

sam (.sam.):

So, \[f(x) = x^{22} - xy + \sin x \cos x - e^x \\ \\ f'(x)= 22x^{21}-(xy'+y)+[\sin(x)(-\sin(x))+\cos(x)\cos(x)]-e^x\]

OpenStudy (error1603):

\[f'(x) = 2x - y + \cos(2x) - exx ( \cos22x - \sin22x = \cos(2x) )\] Using identity. Right?

OpenStudy (error1603):

Using cos22x−sin22x=cos(2x))

sam (.sam.):

Yep

OpenStudy (solomonzelman):

\(\color{black}{f(x,y) = x^{22} - xy + \sin x \cos x - e^x}\) and you want to find \(\color{black}{f_x(x,y)}\)? If this is the question, we treat \(\color{black}{y}\) like a constant (not like a function of \(\color{black}{x}\)). \(\color{black}{f(x,y) = x^{22} - xy + \sin x \cos x - e^x}\) \(\color{black}{f(x,y) = x^{22} - xy + \frac{1}{2}\sin (2x) - e^x}\) I first rewrote the function, and then, \(\color{black}{f_x(x,y) = 22x^{21} -y + \cos (2x) - e^x}\)

sam (.sam.):

\[f(x) = x^{2} - xy + \sin x \cos x - e^x \\ \\ f'(x)= 2x-(xy'+y)+[\sin(x)(-\sin(x))+\cos(x)\cos(x)]-e^x \\ \\ f'(x)=2x-xy'-y+[\cos^2(x)-\sin^2(x)]-e^x \\ \\ f'(x)=2x-xy'-y+\cos(2x)-e^x \]

OpenStudy (error1603):

Thank you! I thought I would never get this.

OpenStudy (solomonzelman):

If this is a partial derivative of a function wrt to \(x\). But, if it is \(\color{black}{y = x^{22} - xy + \sin x\cos x - e^x}\), and you are looking for \(y'\), then you solve it as follows: \(\color{black}{y = x^{22} - xy + \sin x\cos x - e^x}\) \(\color{black}{y = x^{22} - xy + \frac{1}{2}\sin (2x) - e^x}\) \(\color{black}{y' = 22x^{21} - xy' + \cos (2x) - e^x}\) \(\color{black}{y' - xy'= 22x^{21} + \cos (2x) - e^x}\) \(\color{black}{y' (1- x)= 22x^{21} + \cos (2x) - e^x}\) \(\color{black}{\displaystyle y' =\frac{ 22x^{21} + \cos (2x) - e^x}{1-x}}\)

OpenStudy (error1603):

ooooh I see

OpenStudy (error1603):

Thanks

OpenStudy (solomonzelman):

yw